# Spatially-Grounded Document Retrieval via Patch-to-Region Relevance Propagation

### v1

@@ 7.2 Future Directions
Legal Document Retrieval and Citation Grounding. In legal contexts, retrieval precision directly impacts downstream reliability. Recent empirical work demonstrates that even retrieval-augmented legal AI tools hallucinate 17–33% of responses, with errors compounded by coarse retrieval granularity (Magesh et al., 2025). Region-level retrieval with bounding box coordinates provides citation-bounded context—each retrieved region carries verifiable provenance, constraining generation to specific, locatable sources.
@@ 8 Conclusion
We presented a hybrid architecture for spatially-grounded document retrieval that unifies vision-language late interaction with structured OCR extraction. By formalizing the coordinate mapping between ColPali’s patch grid and OCR bounding boxes, we enable region-level retrieval without additional training. Our approach operates entirely at inference time, providing flexibility across OCR systems and VLM backends.
The theoretical analysis establishes precision bounds as a function of patch resolution, derives expected context reduction factors of $5\times$ or more (Theorem 2), and predicts SNR improvements up to $14\times$ under moderate assumptions about score-relevance correlation (Theorem 4). These bounds provide practitioners with guidance on expected performance improvements pending empirical validation.
The two-stage architecture balances computational efficiency with region-level granularity, making the approach practical for large-scale document collections.
We release Snappy as an open-source implementation demonstrating practical applicability. Empirical evaluation is ongoing; we hypothesize that region-level retrieval will improve downstream RAG answer quality while reducing context length and associated costs. Results will be reported in a subsequent revision of this manuscript.
@@ Acknowledgments
The author thanks the ColPali team for their foundational work on vision-language document retrieval and for maintaining an open research ecosystem that enabled this work. The Snappy system implementation is available at https://github.com/athrael-soju/Snappy.

### v2

@@ 7.2 Future Directions
Legal Document Retrieval and Citation Grounding. In legal contexts, retrieval precision directly impacts downstream reliability. Recent empirical work demonstrates that even retrieval-augmented legal AI tools hallucinate 17–33% of responses, with errors compounded by coarse retrieval granularity (Magesh et al., 2025). Region-level retrieval with bounding box coordinates provides citation-bounded context: each retrieved region carries verifiable provenance, constraining generation to specific, locatable sources.
@@ 8 Conclusion
We presented a hybrid architecture for spatially-grounded document retrieval that unifies late-interaction retrieval with structured OCR extraction. By formalizing the coordinate mapping between ColPali’s patch grid and OCR bounding boxes, we enable region-level retrieval without additional training. Our approach operates entirely at inference time, providing flexibility across OCR systems and late-interaction retrieval backends.
On BBox-DocVQA, we evaluate three ColPali-family models. ColQwen3-8B and ColQwen3-4B achieve nearly identical localization accuracy (59.8% and 59.7% hit rate at IoU@0.5, respectively), while ColModernVBERT achieves 45.5%. All models provide substantial token savings compared to full-page retrieval (50.7–58.5%), consistent with the efficiency gains predicted by Theorem˜2.
Category analysis across eight arXiv domains reveals systematic performance variation: computer science and electrical engineering documents achieve higher localization accuracy (Mean IoU $\geq$ 0.65) than economics (0.513) and mathematics (0.382) documents, with physics (0.579–0.590) and quantitative biology (0.616) falling in between.
Analysis of model architecture and scaling effects reveals three empirically distinct regimes. First, ColQwen3-8B and ColQwen3-4B achieve nearly identical performance, indicating that parameter scaling within the same architecture provides negligible localization gains. Second, for architecture-sensitive document types (computer science, electrical engineering), ColQwen3 models substantially outperform ColModernVBERT (up to 35 percentage points at IoU@0.5). Third, for uniformly challenging document types (mathematics), all models converge to similar performance, suggesting fundamental precision limits. This finding has practical implications: ColQwen3-4B offers the best accuracy-efficiency tradeoff, while ColQwen3-8B provides no localization benefit despite doubled parameters.
The two-stage architecture balances computational efficiency with region-level granularity, making the approach practical for large-scale document collections. We release Snappy as an open-source implementation at https://github.com/athrael-soju/Snappy, demonstrating practical applicability for retrieval-augmented generation with reduced context windows and improved precision.
@@ Acknowledgments
The author thanks the ColPali team (ILLUIN Technology) for their foundational work on late-interaction document retrieval and for maintaining an open research ecosystem that enabled this work. We also thank TomoroAI for releasing the ColQwen3 embedding models under Apache 2.0 license, and the Qwen team (Alibaba Cloud) for the Qwen3-VL and Qwen3-Embedding foundation models. The Snappy system implementation is available at https://github.com/athrael-soju/Snappy.